Find the slope and the y-intercept of the line:
3x + 5y + 10 = 0

Answer by Alan3354(69443)

Find the slope and the y-intercept of the line:
3x + 5y + 10 = 0
-------------------
Put the eqn in slope-intercept form, y = mx + b
That means solve for y
y = (-5/3)x - 10/3
Slope m = -5/3
y-int b = -10/3